02 July 2012 UT

D. Gonzalez Huerta
D. Thompson
J. C. Guerra Ramon
O Kuhn

beginning of night start with LUCI authorized on left and IRTC on right (RFBG). Juan Carlos going to test AOS.

23:15 Juan Carlos done with tests. These were successful. Preset and dihtering offsets worked; the AOS displayed all the information correctly. There was a problem that the secondary was NOT offloading. In the PSF DX M2 GUI we see non-zero secondary offload values, however "Sec offload" is No. Juan Carlos traced this to the value set in the *cpp program which was 0.

23:20 Started to make a preset to a bright pointing star to test binorb, but clouds moved in and we closed the dome. I did test this without sending values to the mirrors, but it crashed when it tried a negative square root because the value of M2 Y was -9.0453, so its absolute value was greater than the M2 max value, DXM2max=8.0 set in binorb
ERROR: sqrt of a negative number
  "DXM2Xrange = sqrt( DXM2max*DXM2max - pDXM2Y*pDXM2Y )"
     line 1269: Sandbox$binorb.cl
     called as: `binorb (mode=h)'
But look at the DX M2 position: X=0.569, Y=-9.045, Z=1.367, RX=696.427, RY=583.739. Active Optics offsets for Y/RX in particular are large: Yact=-6.225 and RXact=648.725.

23:20 Closed for clouds.

00:00 Reopened. Clearing corrections. LUCIFER is authorized on LEFT (lfbg) and IRTC on RIGHT (rfbg) Slewing to ACT0232.
newpoint ACT0232 Tel="ACQUIRE" gs=0 BC=1
newpoint ACT0232 Tel="ACQUIRE" gs=0 BC=1

Preset failed, right secondary was RIP. It was already RIP when this was sent and after the testing of Juan Carlos, but he thought it was probably the previous work and it would be OK. David stopped/started IIF, but that did not help. Called Juan Carlos. AOS, AdSec both say M2 is "set" but IIF reports shell as "RIP". David had to exit from all AO software (AOS...etc and IIF) and restart. OK now (00:43).

00:45 Sending another sync preset --- to an M star. Correcting pointing (SX LFBG, LUCI and DX RFBG, IRTC)
  SX DX  
  LFBG RFBG focal station
  LUCI IRTC authorized instrument (though only AGW used)
IE -103 -84  
CA 3 -131  

01:15 Collimation on right was not convering... We saw huge coma on both sides, but left was corrected while on the right, it was not. M2 RY reached limit at ~870" (but lookup table - has RY=818.67 at elev = 63.5 and T1=14.1. Off-axis guide star (1 V=16) for WT10_268.

01:25 Changing to a brighter guide star. Lookup table has :
X=2.58, Y=-3.182, Z=2.723, RX=29.147, RY=818.610

Sending an asynchronous preset on Right only. Collimating, but it does not converge: DX M2 total: X=3.367 and Y=-9.4, Z=1.3, RX=615.7, RY=711.42.
DX M2 AO: X=-1.296, Y=-5.433, Z=0.381, RX=566.16, RY=-135.

LUCIFER is authorized so while we are collimating on right, left M1 is hitting a limit.

Need to do IE/CA to bring star in. Offset from pointing model (calculated):
  SX DX
delta_IE -70.19 -124
delta_CA 3 6

SX_Gregorian.20111018.ptmod on left, LUCI has IE = -32.8097 and CA = 0.00 RFBG_AGw.20111208.ptmod on right, IRTC has IE=40.4619 and CA = -125.7622

Clear the collimation (C00 and active optics) ... did not do ptreset but did set manual pointing offsets are 0.00 Still... to collimate on DX we had to send +2000 of Z7 and Z8. This caused us to hit the limits.

02:05 AGW telemetry crashed.

02:20 Going to a higher elevation star which will bring RY=70 deg. But another AGW problem. AGW #2 probe didn't want to move. Fixed.

~02:40 Closed for clouds.

~02:45 Output from binorb is in /home/lbto/kuhn/2July2012_binorb/ . Data from first test are in 2June2012_binorb. It would be useful to compare output files which list: M1 position (x,y,z,rx,ry,rz), M2 (x,y,z,rx,ry,rz) position, PntOffM1 (x,y,rx,ry) and PntOffM2 (x,y,rx,ry)

2June2012 trial runs
2July2012 (run while collimating on right (only? or both?) but binorb was not sending any corrections... nevertheless the original mirror positions and pntng offsets are recorded).


LBTO@obs5:4 % more jun2coll2_dd_sx.txt
0.8800  0.4091  0.0712 15.1225 -23.446 -0.0142 -2.9331 -1.3642  0.6987 640.331 341.957 -0.0042 
0.3399  0.5501  5.5831 -56.089 -3.8042  0.8705  5.5831 -56.089
14.1820 14.2257 14.1660 -0.0423 69.0306 -1  
0.7659  0.7856  0.0724 46.4526 -13.809 -0.0146 -2.5524 -2.6192  0.6935 673.074 351.910 -0.0164
0.2264  0.9242 36.1365 -46.685 -3.4315 -0.3349 36.1365 -46.685 
14.1820 14.2087 14.1580 -0.0267 68.7579 -1

LBTO@obs4:17 % more jul2a_dd_sx.txt
-0.1403 -0.2659 -0.0004 16.9583 -13.565 -0.0063 -5.9274  1.0440 -0.0949 255.329 -17.439  0.0026
 -0.7095 -0.0281  1.8981 -47.980 -3.1900 -0.1262  1.8981 -47.980 
   14.8373 13.9487 14.6730 -0.0306 62.5483 -1 
-0.1403 -0.2659 -0.0004 16.9586 -13.566 -0.0076 -5.9280  1.0458 -0.0958 255.128 -17.450 -0.0130
 -0.7095 -0.0279  1.8897 -47.980 -3.1900 -0.1256  1.8897 -47.980 
   14.8287 13.8970 14.6690 -0.0195 62.5124 -1

LBTO@obs5:5 % more jun2coll2_dd_dx.txt
-1.0451  1.7074  0.0295 -29.380 16.3745  0.0002  3.4819 -5.6929  1.4802 479.701 659.437 -0.0043 
 -0.2078  1.3397 -5.5743 56.0861  2.7402  1.3574 -5.5743 56.0861 
   13.9720 14.0637 14.1660 -0.0917 69.0289 -1 
-1.0428  1.7168  0.0261 -28.435 16.5986  0.0006  3.4766 -5.7283  1.4674 481.035 656.867 -0.0100
 -0.2630  1.6462 26.9825 66.1830  3.2070 -0.0193 26.9825 66.1830 
   13.9685 14.0487 14.1580 -0.0802 68.7562 -1

LBTO@obs4:16 % more jul2a_dd_dx.txt
-0.1291  0.3640  0.0001 -24.103  9.0956 -0.0217  6.0599 -2.5180  1.1615 28.7510 867.884  0.0157
  0.7095  0.0283 -1.9143 47.9801  3.1900  0.1273 -1.9143 47.9801 
   14.3700 14.2387 14.6820  0.1608 62.5474 -1 
-0.1291  0.3640  0.0001 -24.103  9.0947 -0.0209  6.0598 -2.5180  1.1616 28.7376 867.902 -0.0068 
  0.7095  0.0283 -1.9143 47.9801  3.1900  0.1273 -1.9143 47.9801 
   14.3700 14.2387 14.6820  0.1608 62.5114 -1

03:00 Closed for the night...

  • LBC
    • Turned OFF other systems
    • Turned OFF HK
    • Dave shutdown CMU

  • MODS
    • Took a final set of sieveSnap images
    • Ran modsSleep.pro
    • MODS custom masks removal will be done this morning, between 8-10am by Dave T. & Jeff Urban.
    • ModsShutDown.pro and stopping of MODS services will be done after the custom masks are removed, per Step 1 of MODS Instrument Dismount Preparations (2012 25 June).

  • LUCI
    • No mask in FPU

Notes on binocular range balancing added later on 2 July 2012

  • Question: How could we copoint and collimate on 2 June and not on 2 July
    • In June, used 2012C(?) and now 2012F
    • In July there was an X, Y translation of SX M2 ... Global Offsets in SXSMLUCIFERCollimation.dat are:
      • -1.200 0.500 -1.000 0.000 0.000 0.000 Global Offsets
      • But model was updated only in 2011 December 10, so this LUCIFER model was in place for both.
    • In June we used LBBG and IRTC on SX not LFBG and LUCIFER. Hence we probably used SXSMIRTCBCollimation.dat
      • But this also has the same Global Offsets as the LUCIFER one above. in fact, the collimation models are identical
    • Temperatures are incorporated in the collimation models: both M1 and M2. But these are not too different from June to July.
    • LUCIFER TM Collimation model has non-zero tip/tilt Global Offsets while IRTCF,C,B do not:
# Adjusted focal plane tilt with Tip=+75 JMH/AR/DLM 21-SEP-2010
# Zeroed the global offsets - alignment is in the OSS.cfg file. JMH 30-SEP-2010
# Added Global offsets -300 +50 to reduce AGw3 pupil wobble JMH/DLM 04-MAR-2012
-300.000    50.000     0.000     0.000  Global Offsets

-- OlgaKuhn - 02 Jul 2012
I Attachment Action Size Date Who Comment
MODSInstrumentDismountPrep.pdfpdf MODSInstrumentDismountPrep.pdf manage 335 K 02 Jul 2012 - 10:15 OlgaKuhn MODS Instrument Dismount Preparations (2012 June 25)
Topic revision: r3 - 02 Jul 2012, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback